# Energy-Based Concept Bottleneck Models: Unifying Prediction, Concept Intervention, and Conditional Interpretations

### 25 Jan 2024 v1

@@ 1 Introduction
Concept Bottleneck Models (CBMs) (Koh et al., 2020), as a representative class of models, operate by firstly generating concepts given the input and then using these concepts to predict the final label. The vanilla CBMs often fall short in final prediction accuracy compared to black-box models. This is suboptimal because interpretability is crucial for troubleshooting real problems and should ideally enhance, rather than detract from model performance (Rudin et al., 2022). In order to strike a balance between performance and interpretability, several improvements over the vanilla CBMs have been proposed. For example, Concept Embedding Models (CEMs) (Zarlenga et al., 2022) improve CBMs by including positive and negative semantics, while Post-hoc Concept Bottleneck Models (PCBMs) (Yuksekgonul et al., 2022) make use of residual fitting to compensate for limitations in concept learning. Despite recent advances, existing CBM variants (including CEMs and PCBMs) still suffer from the following key limitations:
@@ 2 Related Work
Concept Bottleneck Models (CBMs), as introduced by Koh et al., date back to earlier work that uses attributes predicted by “attribute” classifiers for downstream tasks such as face verification (Kumar et al., 2009) and animal identification (Lampert et al., 2009). CBMs (Koh et al., 2020) use a convolutional neural network (CNN) as a feature extractor and concept predictor to generate (predict) the “bottleneck” concepts; these concepts are then fed into a label predictor (e.g., a classifier) to predict the final class labels. To better trade-off between accuracy and concept-based interpretability, Concept Embedding Models (CEMs) (Zarlenga et al., 2022) build on CBMs to characterize each concept through a pair of positive and negative concept embeddings. Different from CBMs and CEMs, Post-hoc Concept Bottleneck Models (PCBMs) (Yuksekgonul et al., 2022) abandon the end-to-end training regime and use a post-hoc explanation model with additional residual fitting so that they can use residual features from the input to further improve final accuracy. Probabilistic Concept Bottleneck Models (ProbCBMs) (Kim et al., 2023) incorporate probabilistic embeddings to enable uncertainty estimation of concept prediction. Fully unsupervised ante-hoc concept learning methods (Alvarez Melis & Jaakkola, 2018; Sarkar et al., 2022), utilize a concept encoder to extract concepts along with their corresponding relevances from an image. There are a diverse set of CBM variants (Barbiero et al., 2023; 2022; Havasi et al., 2022; Ghosh et al., 2023a; b; Yang et al., 2023; Sarkar et al., 2022; Oikarinen et al., 2023), each addressing problems from their unique perspectives. This diversity underscores the vitality of research within this field.
@@ 3 Energy-Based Concept Bottleneck Models
    Conditional Interpretations (e.g., $p({\bm{c}}|{\bm{y}})$ or $p(c_{k}|{\bm{y}},c_{k^{\prime}})$). Interpret the model by providing conditional probabilities such as $p(c_{k}|{\bm{y}},c_{k^{\prime}})$ (i.e., given an image with class label ${\bm{y}}$ and concept $c_{k^{\prime}}$, what is the probability that the model correctly predicts concept $c_{k}$).
@@ 4.1 Experiment Setup
Baselines and Implementation Details. We compare our ECBM with state-of-the-art methods, i.e., concept bottleneck model (CBM(Koh et al., 2020), concept embedding model (CEM(Zarlenga et al., 2022), post-hoc concept bottleneck model (PCBM(Yuksekgonul et al., 2022), and probabilistic concept bottleneck model (ProbCBM(Kim et al., 2023). We use ResNet101 (He et al., 2016) as the feature extractor $F$ for all evaluated methods. We use the SGD optimizer during the training process. We use $\lambda_{c}=0.3$ and $\lambda_{g}=0.3$. See Appendix B for more details.
@@ 5 Conclusion and Limitations
In this paper, we go beyond typical concept-based prediction to identify the problems of concept correction and conditional interpretation as valuable tools to provide concept-based interpretations. We propose Energy-based Concept Bottleneck Models (ECBMs), the first general method to unify concept-based prediction, concept correction, and conditional interpretation as conditional probabilities under a joint energy formulation. With ECBM’s unified interface, we derive a set of algorithms to compute different conditional probabilities by composing different energy functions. Limitations of our method include the need for iterative optimization during inference and concept annotation. Generalizing our ECBMs to handle out-of-distribution detection, discover new concepts, adapt label-free methods, and address different downstream tasks would be interesting future work.
@@ C.5 Intervention
We conducted experiments that address the leakage problem as described in  (Havasi et al., 2022). When fewer concepts were given in the training process, the model should not learn as good as the one with all concepts known because of less conceptual information. If the model holds strong performance regardless of concept quantities, it is more likely to suffer from information leakage, i.e., learning concept-irrelevant features directly from image input. This will potentially detriment the accountability of the model interpretation. During training, we provide different quantities of concept groups available for model to learn, then tested the models based on these available concepts. Fig. 11 shows the results. We found that our model performs gradually better when more concept groups are given during training process, instead of intervention-invariant high performances. This indicates that our model suffers less from information leakage, hence providing more reliable interpretations.

### 26 Feb 2024 v2

# Energy-Based Concept Bottleneck Models: Unifying Prediction, Concept Intervention, and Probabilistic Interpretations

@@ 1 Introduction
Concept Bottleneck Models (CBMs) (Koh et al., 2020), as a representative class of models, operate by firstly generating concepts given the input and then using these concepts to predict the final label. The vanilla CBMs often fall short in final prediction accuracy compared to black-box models, creating a potentially unnecessary performance-interpretability trade-off (Rudin et al., 2022). To improve on such trade-off, Concept Embedding Models (CEMs) (Zarlenga et al., 2022) improve CBMs by including positive and negative semantics, while Post-hoc Concept Bottleneck Models (PCBMs) (Yuksekgonul et al., 2022) make use of residual fitting to compensate for limitations in concept learning. Despite recent advances, existing CBM variants (including CEMs and PCBMs) still suffer from the following key limitations:
@@ 2 Related Work
Concept Bottleneck Models (CBMs) (Koh et al., 2020; Kumar et al., 2009; Lampert et al., 2009) use a feature extractor and a concept predictor to generate the “bottleneck” concepts, which are fed into a predictor to predict the final class labels. Concept Embedding Models (CEMs) (Zarlenga et al., 2022) build on CBMs to characterize each concept through a pair of positive and negative concept embeddings. Post-hoc Concept Bottleneck Models (PCBMs) (Yuksekgonul et al., 2022) use a post-hoc explanation model with additional residual fitting to further improve final accuracy. Probabilistic Concept Bottleneck Models (ProbCBMs) (Kim et al., 2023) incorporate probabilistic embeddings to enable uncertainty estimation of concept prediction. There are a diverse set of CBM variants (Barbiero et al., 2023; 2022; Havasi et al., 2022; Ghosh et al., 2023a; b; Yang et al., 2023; Sarkar et al., 2022; Oikarinen et al., 2023), each addressing problems from their unique perspectives. This diversity underscores the vitality of research within this field.
@@
Unsupervised Concept-Based Models, unlike CBMs, aim to extract concepts without concept annotations. This is achieved by introducing inductive bias based on Bayesian deep learning with probabilistic graphical models (Wang et al., 2019; Wang & Yeung, 2016; 2020; Wang & Yan, 2023; Xu et al., 2023), causal structure (Lin et al., 2022), clustering structure (Chen et al., 2019; Ma et al., 2023), or interpretability desiderata (Alvarez Melis & Jaakkola, 2018).
@@ 3 Energy-Based Concept Bottleneck Models
    Conditional Interpretations (Wang et al., 2019) (e.g., $p({\bm{c}}|{\bm{y}})$ or $p(c_{k}|{\bm{y}},c_{k^{\prime}})$). Interpret the model using conditional probabilities such as $p(c_{k}|{\bm{y}},c_{k^{\prime}})$ (i.e., given an image with class label ${\bm{y}}$ and concept $c_{k^{\prime}}$, what is the probability that the model correctly predicts concept $c_{k}$).
@@ 4.1 Experiment Setup
Baselines and Implementation Details. We compare our ECBM with state-of-the-art methods, i.e., concept bottleneck model (CBM(Koh et al., 2020), concept embedding model (CEM(Zarlenga et al., 2022), post-hoc concept bottleneck model (PCBM(Yuksekgonul et al., 2022), and probabilistic concept bottleneck model (ProbCBM(Kim et al., 2023). We use ResNet101 (He et al., 2016) as the feature extractor $F$ for all evaluated methods. We use the SGD optimizer during the training process. We use $\lambda_{c}=0.3$ and $\lambda_{g}=0.3$. For the propositions, we have implemented a hard version (yielding $0/1$ output results) for computing probabilities. See Appendix B for more details.
@@ 5 Conclusion and Limitations
In this paper, we go beyond typical concept-based prediction to identify the problems of concept correction and conditional interpretation as valuable tools to provide concept-based interpretations. We propose ECBM, the first general method to unify concept-based prediction, concept correction, and conditional interpretation as conditional probabilities under a joint energy formulation. Future work may include extending ECBM to handle uncertainty quantification using Bayesian neural networks (Wang & Wang, 2023), enable unsupervised learning of concepts (Ma et al., 2023) via graphical models within the hierarchical Bayesian deep learning framework (Wang & Yeung, 2016; 2020), and enable cross-domain interpretation (Wang et al., 2020; Xu et al., 2022; Liu et al., 2023).
@@ C.5 Intervention
We conducted experiments that address the leakage problem as described in  (Havasi et al., 2022). When fewer concepts were given in the training process, the model should not learn as good as the one with all concepts known because of less conceptual information. If the model holds strong performance regardless of concept quantities, it is more likely to suffer from information leakage, i.e., learning concept-irrelevant features directly from image input. This will potentially detriment the accountability of the model interpretation. During training, we provide different quantities of concept groups available for model to learn, and then test the models based on these available concepts. Fig. 11 shows the results. We found that our model performs gradually better when more concept groups are given during training process, instead of intervention-invariant high performances. This indicates that our model suffers less from information leakage, hence providing more reliable interpretations.